# ViLCo-Bench: VIdeo Language COntinual learning Benchmark

### v1

@@ 1 Introduction
However, existing CL methods are mostly designed for a single modality of data—be it image [24], text [3], audio [43], or video [40, 28]—without considering multiple data modalities and the variety of tasks they entail. Some recent work has explored multimodal data, but only for still-image and textual data [10, 40]. With the rise of embodied AI devices and the abundance of sensor data, there is an urgent need for multimodal ML models to learn collaboratively from diverse data sources. This is crucial for empowering embodied AI agents with natural language understanding while mastering other modalities, such as platforms [17, 25] for human-centric question-answering from videos. Current continual learning benchmarks, focusing mainly on images, videos, or text-image combinations with clean category annotations, may not effectively evaluate continual learning in a video-language setting. Thus, our initial step is to establish a new benchmark in video-language settings, encompassing more challenging multimodal tasks (non-classification tasks).
@@
Our Contributions: Within the realm of multimodal continual learning (CL) tasks, specifically in the video-language domain, there is a notable absence of a unified definition for video-text CL tasks, hindering research progress. Contemporary approaches predominantly emphasize class-incremental CL, which is foundational in the uni-modal domain but may not effectively address the unique challenges of video-language CL tasks. Thus, there is a pressing need for new challenges and benchmarks tailored to the multimodal domain. Despite recent advances in multimodal CL, they primarily focus on text and still images. Existing models for continual learning in streaming data [31] are also limited. Therefore, we propose the creation of a video-text continual learning benchmark to provide a standardized platform for evaluating models in the challenging text-video context.
@@
The main contributions are: 1. We propose the first standardized benchmark in multimodal continual learning for video data, defining protocols for training and metrics for evaluation. This standardized framework allows researchers to effectively compare models, driving advancements in AI systems that can continuously learn from diverse data sources. 2. We define the setup for three recent multimodal tasks in a continual learning setup: Moment Query (MQ), Natural Language Query (NLQ), and Visual Query (VQ). We provide systematic insights into the challenges, gaps, and limitations of each video-text CL tasks. 3. We provide a comparison against four state-of-the-art (SOTA) models in video and video continual learning for each benchmark setup. We prepared a curated dataset suitable for multimodal continual learning tasks using the well-known Ego4D dataset.
@@ 2 Backgrounds and Related Works
Recently, different benchmarks have been introduced for continual learning purposes in different tasks and modalities. CoRE50 [24], Stream-51 [32], are CLeAR [23] are among the most well-known benchmarks for continuous object recognition and streaming classification from still images. vCLIMB is the first that provides a standard benchmark for video continual learning [40]. The vCLIMB benchmark also proposes a novel approach that resolves the complexity of storing videos in memory by employing a temporal consistency loss, which reduces the number of frames stored. However, it focuses on class-incremental classification tasks and does not consider multimodal tasks.
On the other side, [47] provides a visual-question-answering (VQA) benchmark that includes image and text, but it still lacks video data and has not been covered for CL purposes. With recent advancements in multimodal foundation models and the capability of generative models in attending to different modalities (e.g., text, image, audio, video, code, etc.), the application of continual learning in multimodal data is becoming more diverse and inevitable. Despite the importance of the topic, there are no such benchmarks available yet, hence this will be the focus of this work. Table 1 summarizes existing benchmarks.
@@
Recent foundation models aim to learn transferable knowledge shared among modalities (e.g., video and corresponding text) that are generalizable to many downstream tasks [11, 6, 29]. Among these pre-training models, EgoVLP [22, 27] proposed the first comprehensive vision-language dataset using publicly available egocentric video dataset (Ego4D [14]) and formulated applicable tasks for vision-language processing such as Moment Query and Natural Language Query, along with vision only tasks such as instance retrieval and object detection tasks. However, they did not consider continual learning setups in their benchmark. In this work, we introduce major settings for vision-text CL. More reviews on existing CL methods are provided in Appendix A.
@@ 3 ViLCo-Bench: Video-Language Continual Learning Benchmark
Unlike traditional continual learning tasks, our proposed video-language tasks go beyond classification and introduce challenges from cross-modal inference and temporal complexity in videos. We present three unique video-language continual learning tasks to evaluate existing methods and provide a curated dataset with annotations for each episodic memory task. Additionally, we propose novel memory-efficient models tailored for ViLCo-Bench. We design specific evaluation metrics for these tasks and summarize the leaderboard on ViLCo-Bench across various experimental setups.
@@ 3.1 Unique Video-Language Continual Learning Tasks
We designed ViLCo-Bench, a novel continual learning setting, based on the largest egocentric video-language dataset Ego4d [14]. Episodic memory refers to the ability to remember specific first-person experiences, such as "what did I eat" and "who did I sit by on my first flight to France?", which are different from semantic memory tasks like "What’s the capital of France?". This distinction becomes crucial in multimodal continual learning tasks, where models must retain and recall specific episodes in the past. In contrast to existing continual learning benchmarks, ViLCo-Bench presents three unique challenges. (1) Exploring the Video-Language Domain: ViLCo-Bench is the first benchmark to delve into the video-language domain in continual learning. The videos in this benchmark are significantly longer than those in previous video-based benchmarks. The average video length in our benchmark is 10 minutes, which is 85 times longer than the average video length in UCF101 [35]. (2) Beyond Classification Tasks: traditional continual learning benchmarks primarily focus on classification tasks. However, ViLCo-Bench introduces novel video-language tasks that require models to understand the semantic relationships between language queries and video content. (3) Complex Annotation Splitting: The annotation in video-language tasks (i.e., extracting temporal windows that each object/activity has appeared on the screen), is challenging. Moreover, there is overlap across labels which changes the problem to a multi-label multimodal learning problem. To address the aforementioned challenges, we introduce a novel setting called Query-Incremental Learning (QIL). The QIL includes three task formulations, each introducing unique complexities and necessitating specific model adaptations. As illustrated in Fig. 3, these tasks include:
Moments Queries (MQ): For each MQ sub-task, the model is tasked with recognizing and recalling all temporal windows corresponding to queries about different human activities or specific moments. These queries can include prompts like "When did I drink a beverage?" or "When did I use the phone?" Our task-splitting strategy is based on the action categories within the query space.
Natural Language Queries (NLQ): In NLQ tasks, the model is required to interpret and respond to natural language queries related to specific episodic memories. For instance, given a query like "What restaurant did I visit last Saturday?", the model needs to retrieve and identify the relevant temporal window from its episodic memory.
Visual Queries (VQ): VQ tasks require the processing of visual information linked to specific episodic memories. For instance, consider a scenario where the model is presented with an image of a crowded airport; it should be capable of recalling detailed localizations, such as answering the query "Where is the toolbox when I repair the bicycle?" using visual cues from the image. The division of VQ tasks is structured around object categories.
@@ 3.2 Dataset Curation
We used Ego4d dataset [14] consisting of 3670 hours of first-person videos for episodic memory tasks based on egocentric videos. The tasks are to predict temporal windows or 2-dimensional bounding boxes from various queries, including Moments Queries (MQ), Natural Language Queries (NLQ), and Visual Queries (VQ). However, applying the traditional class-incremental setting directly is not suitable for video-language continual learning due to the complexity of multimodal tasks and the varying label formats Therefore, we select videos and their corresponding queries to create three subsets for continual learning: ViLCo-Bench-MQ, ViLCo-Bench-NLQ, and ViLCo-Bench-VQ.
In ViLCo-Bench-MQ, we sample 165 hours of videos consisting of $110$ action categories that includes over $10,000$ videos with the duration varying between 8 to 10 minutes. We divide the task into five sub-tasks, each containing $22$ action category queries. Annotations for each video may involve multiple classes and timestamps. ViLCo-Bench-NLQ consists of 136 hours of video. However, ViLCo-Bench-NLQ is more challenging because this subset leverages natural language as queries, such as "Where is Object X?", where "X" could be arbitrary textual descriptions. We summarize 13 query templates resulting in 13 sub-tasks. Query templates for NLQ scenarios are provided in Appendix B. ViLCo-Bench-VQ includes of 167 hours of videos based on the object categories from image queries. we divided it into 5 splits each containing over 400 categories.
When partitioning existing tasks into non-overlapping sub-tasks for CL, we consider the long-tail distribution of queries. Specifically, we strive to balance the frequency of different queries during the partitioning process to ensure equitable exposure to various types of queries. However, a significant challenge arises as over $90\%$ of the videos involve multiple classes spanning different subsets. To address this, we ensure that each video is exclusive to one subset by prioritizing queries of higher frequency and excluding queries that appear across multiple subsets. This method of allocation minimizes overlap and enhances the clarity of learning objectives within each subset. This approach facilitates more effective training and evaluation of models under continual learning scenarios.
@@ 3.3 ViLCo Model
We investigate the factors hindering existing continual learning methods [19, 1] within video-language continue learning setting. Most of these techniques are developed from uni-modal benchmarks and focus on classification tasks. ViLCo-Bench introduces long-term videos and flexible natural language descriptions, as well as various multimodal tasks. To resolve the misalignment between video and language within multimodal CL, we propose a memory-efficient framework by adopting self-supervised learning.The overall architecture of ViLCo-Bench is shown in Fig 4 and consists of five parts: video feature encoder, text encoder, episodic memory module, cross-modal encoder and task-specific heads. Specifically, given a video clip and corresponding language query, first, visual and textual embeddings are exploited using the video feature and textual encoders. For each sub-task in continual learning, we update the episodic memory with a few new samples to mix the previous tasks with the current data. Then, visual and textual features are fed into the cross-modal encoder to obtain fused representations at multiple levels. Finally, the cross-modal encoder fuses features to generate predictions for various tasks (e.g., MQ, NLQ and VQ).
@@
Figure 4: Overview of ViLCo-Bench employing episodic memory modules.
Video Feature Encoder. In NLQ and MQ tasks, we adopt EgoVLPv2 [27] as the visual encoder. For VQ tasks, the visual encoder is from a DINO pre-trained ViT [46]. We freeze the video feature encoder and independently extract visual features for sampled video frames. We follow the same frame sampling strategy as EgoVLPv2.
Textual Encoder. We employ a CLIP (ViT-L/14) [29] to exploit text embeddings. As the MQ task includes multiple action categories in queries, we concatenate them together.
@@
Cross-modal Encoder. We fuse video and language features with a Transformer network. Inspired by [21, 34], we also adopt a feature pyramid network to improve spatial-temporal representations by extracting features at multiple levels. This helps to understand temporal information in fine detail and precise localization prediction.
@@
Task-specific Head. Given the variations in targets and annotations across different tasks, such as natural language query (NLQ), we have implemented task-specific heads to cater to these distinctions. For moments query (MQ) and NLQ, we utilize a classification head and a localization head, respectively, to produce dense temporal predictions. It is important to note that in NLQ, the number of classes is typically one, as the queried events generally occur only once within an entire video. For visual query (VQ), our model includes a head designed to estimate the probability of the anchor-level queried object and a regression head to precisely adjust the coordinates of the bounding boxes.
@@
Episodic Memory Module. As ViLCo-Bench introduces tasks involving long-term video analysis and multimodal interactions, it necessitates a paradigm shift towards training a more effective memory system. Traditional methods rely on rehearsal (short-term) buffers, which store a subset of seen examples from previous tasks. These stored examples are periodically rehearsed or re-introduced into the learning process when new data is being learned. One of the updating strategies is to randomly select a set of examples from the incoming data stream, either replacing older entries or filling in available memory space. In video language continual learning, maintaining most frame-level features in rehearsal buffers occupies substantial storage space, making it inefficient for long-term episode analysis and multimodal learning. Thus, we propose a long-term replay buffer designed to dynamically facilitate episodic memory recall across varied events and tasks, as shown in Fig. 4.
@@
The primary objective of the long-term replay buffer is to preserve historical knowledge across similar queries from distinct tasks, thereby mitigating the risks of significant memory degradation. The buffer is characterized by a compact set of learnable parameters, which occupy a constrained memory space, formally defined as $\{(\mathcal{P}_{1},\mathcal{K}_{1}),(\mathcal{P}_{2},\mathcal{K}_{2}),...,(\mathcal{P}_{m},\mathcal{K}_{m})\}$ where $\mathcal{P}$ represents the encoded prompt for episodic memory. $\mathcal{K}$ is a learnable key associated with value $\mathcal{P}$, dimensioned at $L\times D$. $L$ denotes the length of inputs, and $D$ indicates the embedding size, while $m$ represents the total memory capacity. We would like to automatically select the prompt by matching the language query and key matrix. To this end, we calculate the similarity between query embeddings and prompt keys, selecting the top-N prompts that best represent episodic memory. We score the similarity between query embeddings and prompt keys to select the top-N prompts as the representations for episodic memory. These selected prompt embeddings are then used to replace the original query embeddings, making the framework task-agnostic and revisiting analogous knowledge across different tasks. At each training step, we update $K$ and $P$ by minimizing the end-to-end training loss function.
@@
Self-supervised Learning with Narrations. We employ SSL to improve the cross-modal representations by incorporating augmented versions of video narrations. The difficulty of collecting annotations of video-language tasks hinders the generalization of models. Inspired by self-supervised works [29, 9], we leverage additional video narrations provided by Ego4d dataset to reduce the misalignment between videos and language. Video narrations provide descriptive information for each video clip, and we adopt contrastive learning to learn robust representations. To avoid forgetting issues in CL, we sample and store a few narration examples from previous tasks to construct the negative pairs in the short-term memory (See Appendix C).

### v2

@@ 1 Introduction
However, existing CL methods are mostly designed for a single modality of data—be it image [27, 31], text [4], audio [49], or video [46]—without considering multiple data modalities and the variety of tasks they entail. Some recent work has explored multimodal data, but only for still-image and textual data [12, 46]. With the rise of embodied AI devices and the abundance of sensor data, there is an urgent need for multimodal ML models to learn collaboratively from diverse data sources. This is crucial for empowering embodied AI agents with natural language understanding while mastering other modalities, such as platforms [20, 28] for human-centric question-answering from videos. Current continual learning benchmarks, focusing mainly on images, videos, or text-image combinations with clean category annotations, may not effectively evaluate continual learning in a video-language setting. Thus, our initial step is to establish a new benchmark in video-language settings, encompassing more challenging multimodal tasks (non-classification tasks).
@@
Our Contributions: Within the realm of multimodal continual learning (CL) tasks, specifically in the video-language domain, there is a notable absence of a unified definition for video-text CL tasks, hindering research progress. Contemporary approaches predominantly emphasize class-incremental CL, which is foundational in the uni-modal domain [46, 36, 45] but they do not address the unique challenges of video-language CL tasks [54]. Thus, there is a pressing need for new challenges and benchmarks tailored to the multimodal domain. Despite recent advances in multimodal CL, they primarily focus on text and still images. Existing models for continual learning in streaming data [36] are also limited. Therefore, we propose the creation of a video-text continual learning benchmark to provide a standardized platform for evaluating models in the challenging text-video context.
@@
The main contributions are: 1. We propose the first standardized benchmark in multimodal continual learning for video data, defining protocols for training and metrics for evaluation. This standardized framework allows researchers to effectively compare models, driving advancements in AI systems that can continuously learn from diverse data sources. 2. We define the setup for three recent multimodal tasks in a continual learning setup: Moment Query (MQ), Natural Language Query (NLQ), and Visual Query (VQ). We provide systematic insights into the challenges, gaps, and limitations of each video-text CL tasks. 3. We provide a comparison against four state-of-the-art (SOTA) models in video and text continual learning for each benchmark setup. We prepared a curated dataset suitable for multimodal continual learning tasks using the well-known Ego4D dataset.
@@ 2 Backgrounds and Related Works
Recently, different benchmarks have been introduced for continual learning purposes in different tasks and modalities. CoRE50 [27], Stream-51 [37], and CLeAR [26] are among the most well-known benchmarks for continuous object recognition and streaming classification from still images. vCLIMB is the first that provides a standard benchmark for video continual learning [46]. The vCLIMB benchmark also proposes a novel approach that resolves the complexity of storing videos in memory by employing a temporal consistency loss, which reduces the number of frames stored. However, it focuses on class-incremental classification tasks and does not consider multimodal tasks.
@@
On the other side, [55] provides a visual-question-answering (VQA) benchmark that includes image and text, but it still lacks video data and has not been covered for CL purposes. With recent advancements in multimodal foundation models and the capability of generative models to attend to different modalities (e.g., text, image, audio, video, code, etc.), the application of continual learning in multimodal data is becoming more diverse and inevitable. Despite the importance of the topic, there are no such benchmarks available yet, hence this will be the focus of this work. Table 1 summarizes existing benchmarks.
@@
Recent foundation models aim to learn transferable knowledge shared among modalities (e.g., video and corresponding text) that are generalizable to many downstream tasks [13, 7, 32]. Among these pre-training models, EgoVLP [25, 30] proposed the first comprehensive vision-language dataset using publicly available egocentric video dataset (Ego4D [16]) and formulated applicable tasks for vision-language processing such as Moment Query and Natural Language Query, along with vision only tasks such as instance retrieval and object detection tasks. However, they did not consider continual learning setups in their benchmark.More reviews on existing CL methods are provided in Appendix A.
@@ 3 ViLCo-Bench: Video-Language Continual Learning Benchmark
Unlike traditional continual learning tasks, ViLCo-Bench goes beyond classification and introduces challenges of cross-modal inference and temporal complexity in videos. We present three unique video-language continual learning tasks to evaluate existing methods and provide a curated dataset with annotations for each episodic memory task. Additionally, we propose novel memory-efficient models tailored for ViLCo-Bench.
@@ 3.1 Unique Video-Language Continual Learning Tasks
ViLCo-Benchproposes a novel continual learning setting, based on the largest egocentric video-language dataset Ego4D [16] specifically focusing on episodic memory tasks. Episodic memory refers to the ability to remember specific experiences and make past videos queriable [33], such as "what did I eat" and "who did I sit by on my first flight to France?", which are different from semantic memory tasks like "What’s the capital of France?". This distinction becomes crucial in multimodal continual learning tasks, where models must retain and recall specific episodes in the past. In contrast to existing continual learning benchmarks, ViLCo-Bench presents three unique challenges. (1) Exploring the Video-Language Domain: ViLCo-Bench is the first benchmark to delve into the video-language domain in continual learning. The videos in this benchmark are significantly longer than those in previous video-based benchmarks. The average video length in our benchmark is 10 minutes, which is 85 times longer than the average video length in UCF101 [41]. (2) Beyond Classification Tasks: traditional continual learning benchmarks primarily focus on classification tasks. However, ViLCo-Bench introduces novel video-language tasks that require models to understand the semantic relationships between language queries and video content. (3) Complex Annotation Splitting: The annotation in video-language tasks (i.e., extracting temporal windows that each object/activity has appeared on the screen), is challenging. Moreover, there is overlap across labels which changes the problem to a multi-label multimodal learning problem. To address the aforementioned challenges, we introduce a novel setting called Query-Incremental Learning (QIL). The QIL includes three task formulations, each introducing unique complexities and necessitating specific model adaptations. As illustrated in Fig. 3, these tasks include:
Moments Queries (MQ): For each MQ sub-task, the model identifies and recalls temporal windows for queries about specific human activities, such as "When did I drink a beverage?" or "When did I use the phone?" We split tasks based on action categories within the query space.
Natural Language Queries (NLQ): In NLQ tasks, the model interprets and responds to queries about specific episodic memories. For example, given "What restaurant did I visit last Saturday?" the model must retrieve and identify the relevant temporal window.
Visual Queries (VQ): VQ tasks involve processing visual information tied to episodic memories. For example, given a video of a garage workshop, the model should recall specific details like the location of a toolbox when asked, "Where is the toolbox when I repair the bicycle?" VQ tasks are divided based on object categories.
@@ 3.2 Dataset Curation
We used Ego4D dataset [16] consisting of 3670 hours of egocentric videos . The tasks are to predict temporal windows or 2-dimensional bounding boxes from various queries mentioned before. These multimodal tasks introduce new challenges compared to unimodal (vision-only) models such as noisy videos due to quick head movements and limited field of view, free-form queries in natural language, and tiny response windows within lengthy video footage [33]. Therefore, we select videos and their corresponding queries and narrations to create three subsets for continual learning: ViLCo-Bench-MQ, ViLCo-Bench-NLQ, and ViLCo-Bench-VQ.
In ViLCo-Bench-MQ, we sample 165 hours of videos consisting of $110$ action categories that include over $10,000$ videos varying between 8 to 10 minutes. We divide the task into five sub-tasks, each containing $22$ action category queries. Annotations for each video may involve multiple classes and timestamps. ViLCo-Bench-NLQ consists of 136 hours of video. However, ViLCo-Bench-NLQ is more challenging because this subset leverages natural language as queries, such as "Where is Object X?", where "X" could be arbitrary textual descriptions. We summarize 13 query templates resulting in 13 sub-tasks. Query templates for NLQ scenarios are provided in Appendix B. ViLCo-Bench-VQ consists of 167 hours of videos based on the object categories from image queries. we divided it into 5 splits each containing over 400 categories.
To avoid the label overlapping issue across tasks and samples [40], we adopted a partitioning strategy to create and split training sets for different tasks and sub-tasks. Since over 90% of videos contain multiple classes spanning different subsets, we ensured each video was assigned to only one subset. We did this by prioritizing higher-frequency queries and excluding those that appeared in multiple subsets, thereby minimizing overlap and enhancing the clarity of learning objectives. This approach supports more effective training and evaluation of models in continual learning scenarios.
@@ 3.3 ViLCo Model
We investigate the impact of factors developed by existing continual learning methods [22, 1] within the context of video-language continual learning setting. Most of these techniques are developed for uni-modal benchmarks and focus on classification tasks. ViLCo-Bench introduces long-term videos and flexible natural language descriptions across various multimodal tasks. To align the video and language modalities within multimodal CL, we propose a memory-efficient framework by adapting a self-supervised learning pre-training mechanism and incorporating short-term and long-term memory modules (replay buffer). Fig 4 shows the overall architecture of ViLCo-Bench which consists of four key modules as introduced below:
@@
Figure 4: Overview of ViLCo-Bench employing episodic memory modules. All the modules except task-specific headers are shared among tasks and continually get updated.
Video and Text Encoder. Regarding the video data, in NLQ and MQ tasks, we adopt EgoVLPv2 [30] as the visual encoder. For VQ tasks, we used pre-trained ViT visual encoder by DINO [53]. We follow the same frame sampling strategy as EgoVLPv2. Moreover, we employ a CLIP (ViT-L/14) [32] to extract text embeddings across all types of tasks. In MQ tasks with multiple action categories per query, we concatenate them together.
@@
Cross-modal Encoder. To fuse the video and language features we employ a Transformer network. Inspired by [24, 39], we also adopt a feature pyramid network to improve spatial-temporal representations by extracting features at multiple levels. This helps to understand temporal information in fine detail and precise localization prediction.
@@
Task-specific Head. Given the variations in targets and annotations across different tasks, such as NLQ, we have implemented task-specific heads to address these distinctions. For MQ and NLQ tasks, we use a classification head and a localization head, respectively, to generate dense temporal predictions. It is important to note that in NLQ, the number of classes is typically one, as the queried events generally occur only once within an entire video. For VQ, our model employs a head designed to estimate the probability of the anchor-level queried object and a regression head to precisely adjust the coordinates of the bounding boxes.
@@
Episodic Memory Module. To address long-term video analysis and multimodal interactions, it requires a more effective memory system. Existing CL models mostly rely on rehearsal (short-term) buffers, which store a subset of seen examples from previous tasks [52, 17, 35]. These stored examples are periodically rehearsed or re-introduced into the learning process when new data is being learned. One of the updating strategies is to randomly select a set of examples from the incoming data stream, either replacing older entries or filling in available memory space [46]. In video language continual learning, maintaining most frame-level features in rehearsal buffers occupies substantial storage space, making it inefficient for long-term episode analysis and multimodal learning [34]. Thus, we propose a dual-memory module consisting of short-term and long-term replay buffers to dynamically facilitate episodic memory recall across varied events and tasks, as shown in Fig. 4. Both memory modules are initialized using the standard normal distribution.
@@
Short-term memory component stores the text and vision embeddings across the current task. We extract negative samples from short-memory to be used for self-supervised training of the vision and text encoders. The primary objective of the long-term replay buffer is to preserve historical knowledge across similar queries from distinct tasks, thereby mitigating the risks of significant memory degradation. The buffer is characterized by a compact set of learnable parameters$\{(\mathcal{P}_{1},\mathcal{K}_{1}),(\mathcal{P}_{2},\mathcal{K}_{2}),...,(\mathcal{P}_{m},\mathcal{K}_{m})\}$ where $\mathcal{P}$ represents the encoded prompt for episodic memory. $\mathcal{K}$ is a learnable key associated with value $\mathcal{P}$, dimensioned at $L\times D$. $L$ denotes the length of inputs, and $D$ indicates the embedding size, while $m$ represents the total memory capacity. We would like to automatically select the prompt by matching the language query and key matrix. To this end, we calculate the similarity between query embeddings and prompt keys, selecting the top-N prompts that best represent episodic memory. These selected prompt embeddings are then used to replace the original query embeddings, making the framework task-agnostic and revisiting analogous knowledge across different tasks. At each training step, we update $K$ and $P$ by minimizing the end-to-end training loss function. We employ a contrastive loss to regulate and update the prompt pool. This loss helps select prompts that minimize the distance between the task embeddings and the relevant keys, while maximizing the distance from irrelevant prompts (See Appendix LABEL:apx:long_term_buffer).
@@
Self-supervised Learning with Narrations. We use SSL to enhance cross-modal representations by incorporating augmented video narrations. The challenge of collecting annotations for video-language tasks limits model generalization. Inspired by self-supervised methods [32, 11], we leverage additional video narrations from the Ego4d dataset to reduce video-language misalignment. By adopting contrastive learning, we learn robust representations from these narrations. To prevent forgetting in CL, we sample and store a few narration examples from previous tasks to construct negative pairs in short-term memory.(See Appendix C).